• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 01:45:27 on October 31, 2008

    Entry number 245644

    keyword=summary of target progress on Thursday (Target Polarization around 50%)

    One good news today (Thursday) is the finding that the target polarization actually has been reasonably high (~50%). The earlier NMR measurement had a wrong cable connection (NMR preamp connected to 650 Ohm instead of 50 ohm), which caused the polarization number to be off by a large factor (~4).

    Last night, Chiranjib found that the EPR RF amplifier was cold. This afternoon, when replacing the EPR RF amplifier and checking it out, Ye found that the NMR preamp connection was wrong and after correct it, polarization went to ~50%. EPR was still not working even with the new RF amplifier. Jin/Chiranjib checked D2 light last night and found it had plenty. This evening the work is focusing on increasing RF for EPR. Ye and Ye' changed cable length to make it closer to the condition in the target lab (when EPR worked well). Also checked Modulation and it's fine. Jin/Chiranjib is working on laser alignment for the transverse line. Once the alignement is complete (in next hour), will test EPR again. If it does not work, will go into the oven to lower the RF coil to further increase RF power, which will require going to target enclosure. We have got RADCON clearance already for this work.